OpenStudy (anonymous):

Suppose f varies inversely with g and that f=28 when g=2. What is the value of f when g=7?

OpenStudy (mertsj):

f=kg is direct variation f=k/g is inverse variation.

OpenStudy (anonymous):

so i divide 7/2?

OpenStudy (mertsj):

You first plug in the given values for f and g and find k

OpenStudy (anonymous):

k= 28 * 2= 56

OpenStudy (anonymous):

?

OpenStudy (mertsj):

Yes. So now we know that f = 56/g

OpenStudy (mertsj):

So now you can find f when g = 7

OpenStudy (anonymous):

oh ok thanks i have one more question

OpenStudy (mertsj):

Are you going to tell me what it is or should I try and guess?

OpenStudy (anonymous):

Use the given graph of the function f to find the following values for f^(-1) f−1(−4)= 2. f−1(−3)= 3. f−1(0)= 4. f−1(2)= 5. f−1(4)=

OpenStudy (anonymous):

the ans to the previous question is 8

OpenStudy (anonymous):

56/7=8

OpenStudy (mertsj):

yep. good job

OpenStudy (anonymous):

how do i find the inverse of the graph and then look up the f(x)

OpenStudy (mertsj):

Do you see that (2,-4) is on that graph?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

In the inverse function all the ordered pairs are reversed. So (-4,2) is an ordered pair of f^(-1)(x)

OpenStudy (mertsj):

Or, in other words: \[f ^{-1}(-4)=2\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then f^-1 (3)=1

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

what about 0 and 2

OpenStudy (anonymous):

wait is f^-1(2)=0

OpenStudy (mertsj):

It looks to me like f(-1.5)=0 Would you agree?

OpenStudy (mertsj):

So the function f contains the ordered pair (-1.5,0) So \[f ^{-1}(0)=-1.5\]

OpenStudy (anonymous):

yes i agree with -1.5=0...i was wrong about f^-1(2)=0

OpenStudy (mertsj):

On the graph is the ordered pair (-2.5,2) so \[f ^{-1}(2)=-2.5\]

OpenStudy (anonymous):

would you say f^-1(4)= -4

OpenStudy (mertsj):

Is (-4,4) on the graph you posted?

OpenStudy (anonymous):

nvm, i miscounted

OpenStudy (anonymous):

-3.5,4 is there though right?

OpenStudy (mertsj):

Yes.

OpenStudy (anonymous):

thanks for the help

OpenStudy (mertsj):

yw